# Fractional phase jumps in stochastic systems with tilted periodic double-well potentials

### v1

@@ III.1 Strong damping
We start our analysis with the strong damping case. The main reason is that in this regime the phase dynamics is much simpler than for the intermediate and weak damping. Nevertheless, it is still far from trivial.
III.1.1 Fano factor and phase jumps
@@
Figure 3: (a) Examples of typical low-temperature Fano factor $F=S/2\pi\langle v\rangle$ for strong damping plotted as a function of bias force for different parameters $\alpha$. The inset illustrates how the $0.5$ Fano factor plateau transits into Fano plateau with value one as $\alpha$ increase. (b) A comparison of the $\alpha$-dependence of the single-phase-jumps rates evaluated numerically (solid lines) with the analytical Kramers formula result for the overdamped case [77, 69] (dashed lines). The green bullets represent rates obtained via the simplified model where only $2\pi n$ PJ were taken into account.
@@
In Fig. 3(a) we show the Fano factor in the strongly damped regime represented by $\gamma=5$ for three values of $\alpha$ at temperature $\Theta=0.08$. The Fano factor exhibits a characteristic divergence close to the equilibrium ($i_{b}=0$) due to the finite thermal noise. For $\alpha=0$, it follows formula $F=\frac{1}{2}\coth(\pi i_{b}/2\Theta)$ [69, Eq. (28)] for small $i_{b}$ and $\alpha=0$ with a plateau at the Poissonian value of $F=0.5$ (note that we still normalize $F$ to $2\pi$). This is in agreement with the dominant contribution to the noise from the single PJ [illustrated in Fig. 1(c)] over a distance $d=\pi$ between the equivalent neighboring minima. The situation for the finite $\alpha$ is more complicated. The range in which single PJs are the dominant source of the overall noise is not marked by a clear plateau. Rather for $0<\alpha\ll 1$ [$\alpha=0.1$ in Fig. 3(a)] we observe a slight slope in the Fano factor. This reflects the fact that for finite $\alpha$ the prevailing velocity noise contribution consists of a nontrivial combination of the single phase jumps forward over the distance of $\varphi^{B}_{m}-\varphi^{A}_{m}=d$ and backward by $\varphi^{A}_{m+1}-\varphi^{B}_{m}=2\pi-d$ [see Fig. 1(c)] where $d$ depends both on $\alpha$ and $i_{b}$. As already stated, because $d\neq\pi$ we refer to these events as the fractional phase jumps.
@@
With increasing $\alpha$ the Fano factor increases smoothly in this region from $0.5$ signaling only PJs over the distance $\pi$ to $1$ suggesting a single PJs over $2\pi$ as shown in the inset of Fig. 3(a). Interestingly, for $\Theta=0.08$ the Fano factor approaches one even for values of $i_{b}$ that are still well below $i_{cA}$, therefore, in a regime where both minima still exist. For example, the critical current $i_{cA}$ is approximately $0.396$ at $\alpha=1$, yet its Fano factor follows the analytical result $F=\coth(\pi i_{b}/\Theta)$, valid for systems with only the first harmonic, even below this value. Nevertheless, this can be understood as a consequence of the large $\gamma$ and can be explained using the FCS method together with a simplified model of the elementary PJ processes.
@@
Because only jumps over a single (uneven) maxima are realized in the locked state of the strong-damping case the overall dynamics of this regime can be described using just four rates: $\Gamma^{A}_{1}$ and $\Gamma^{B}_{1}$ for the forward single jumps and $\Gamma^{A}_{-1}$ and $\Gamma^{B}_{-1}$ for the backward single jumps. Moreover, the backward rates can be neglected if the bias force is strong enough. The typical dependencies of $\Gamma^{A}_{1}$ and $\Gamma^{B}_{1}$ on $\alpha$ in this regime represented by the bias force $i_{b}=0.3$ and the damping $\gamma=5$ are plotted in the Fig. 3(b) (solid lines). Both $\Gamma^{A}_{1}$ and $\Gamma^{B}_{1}$ closely follow the Kramers formula for escape across the adjacent barrier for overdamped case $\Gamma^{X}_{1}=\frac{1}{2\pi}\sqrt{|U^{\prime\prime}(\varphi^{x})|U^{\prime\prime}(\varphi^{X})}e^{-\Delta U_{X}/\Theta}$ [77, 69] (where $\Delta U_{X}=U(\varphi^{x})-U(\varphi^{X})$, and $x=a,b$; $X=A,B$) plotted with dashed lines of adequate colors. Note that because of the increasing difference $\Delta U_{B}-\Delta U_{A}$ the ratio
@@
increases exponentially with $\alpha$. Consequently, for high enough $\alpha$ and low enough $\Theta$ the average waiting time for the escape from minimum $A$ $\tau_{A\rightarrow B}=1/\Gamma^{A}_{1}$ is negligible compared to the waiting time for the escape from minimum $B$ ($\tau_{B\rightarrow A}=1/\Gamma^{B}_{1}$). Therefore, in the long term, every single phase jump $B_{m-1}\rightarrow A_{m}$ is immediately followed by a single phase jump $A_{m}\rightarrow B_{m}$. The combination of these two fractional PJs is effectively a complete $2\pi$ phase jump. This is shown in Fig. 3(b) by the green bullets that were obtained via a simplified model where only the $2\pi n$ phase jumps were considered [35]
@@
Their match with the $\Gamma^{B}_{1}$ rate for high enough $\alpha$ is consistent with the Fano-factor value of one in the inset of Fig. 3(a). This has interesting physical consequences. If the temperature is low enough, then, because $\Theta$ is in the denominator of the exponent of Eq. (27), any strongly damped $\varphi$-junction or another equivalent system describable by a double harmonic potential with a finite bias $i_{b}$ will in the steady state resemble a simple single harmonic system. As such, it can be described by the analytical formulas derived for the single-harmonic potential.
@@ III.2 Intermediate and weak damping
The stochastic dynamics of the particle in the tilted double-well periodic potential in the regime of intermediate and weak damping is significantly richer than in the strong damping case. For example, for weak bias force, there are phase jumps over multiple maxima, and for stronger bias, a complicated switching between running and locked solutions is the prevailing source of the velocity noise. Even the retrapping of a particle from the running to locked regime has complex dynamics, as it is discussed below.
@@
We support this claim in Fig. 4(b) where an example of the stationary distribution function $W(v,\varphi)$ is plotted for the parameters (see figure description) close to the maximum of the peak. Here, the two distinct peaks at $v\approx 0$ centered around potential minima signal the two locked states. The continuous ridge that spreads above them is the running phase. In this regime, the prevailing contribution to the overall velocity noise comes from the switching between these three well-separated metastable states.
When we lower $i_{b}$ to the regime in between the thermal divergence at $i_{b}\to 0$ and the switching maximum, multiple fractional phase jumps (MPJ) become the prevailing source of the velocity noise. We now analyze the regimes of phase jumps and switching separately and then show how they transition smoothly into each other.
Figure 4: (a) Three examples of a typical low-temperature Fano factor $F=S/2\pi v$ dependence on the bias force for weak damping. The lines were calculated by the MCF method, and the bullets show the ME results from Eq. (26). Vertical arrows mark the related noiseless retrapping forces. (b) An illustration of the stationary distribution function $W(v,\varphi)$ in the switching regime calculated for $\gamma=0.2$, $\alpha=0.02$, $i_{b}=0.38$, and $\Theta=0.08$. The arrows illustrate the processes and their rates as used in Eq. (39). Namely, the red arrows represent escape from the locked states to the running one, the black arrows represent retrapping of the particle, and the blue arrows are the phase jumps between the locked states $A$ and $B$. The gray dashed curves show the separatrices between the running solution and the locked solution in the well $A$ (solid line) and/or well $B$ (dashed line), respectively.
III.2.1 Phase-jumps
Figure 5: The rates of MPJ (of order $n$) for different values of the bias force $i_{b}$ and ratio $\alpha$. The inset in (a) is a verification of the detailed balance condition. The blue columns describe jumps starting at minimum $A$, red at minimum $B$ (a small horizontal shift is introduced for visibility), and the green columns in panel (c) show the rates obtained via the simplified model where only $2\pi n$ PJ were considered (analogously to Fig. 3(b)).
@@ III.2.1 Phase-jumps
The profile of the rates for potentials with finite $\alpha$ is more complicated. The $n$ dependencies of the PJs rates differ qualitatively between jumps starting in different minima, as well as between odd and even $n$’s. We show this in Figs. 5(b)-(c). The rates $\Gamma_{A\,n}$ and $\Gamma_{B\,n}$ differ by several orders of magnitude for odd $n$’s [see $\alpha=0.2$ case in Fig. 5(c)], but are comparable for even $n$’s. This can be rationalized by analyzing the “trajectories” of the particular PJs following the illustrations in Fig. 1(b)(c).
The jumps $A_{m}\rightarrow A_{m+n/2}$ and $B_{m}\rightarrow B_{m+n/2}$, where $n$ is an even integer, are indeed comparable. Here, the particle had to overcome the same number of maxima $a$ as well as $b$, namely $n/2$. In the idealized case, they traveled the same distance $\pi n$. Therefore, also the related rates are equivalent. However, this is not true for odd jumps representing fractional PJs between minima of different kinds. In the case of an odd $n$ the particle overcomes $(n+1)/2$ of the lower maxima but only $(n-1)/2$ of the higher ones during a $\Gamma^{A}_{n}$ process. The opposite is true for the $\Gamma^{B}_{n}$ process. Furthermore, travel distances differ by $2\pi-2d$. Consequently, the “odd” rates differ significantly ($\Gamma^{A}_{n}\gg\Gamma^{B}_{n}$ ).
@@
Before moving to the switching regime, it is worth stressing that the FCS method works well up to surprisingly high Fano factors ($F\sim 10^{2}$). This is shown in Fig. 4(a) where the bullets represent the Fano factor obtained directly from the rates by Eq. (26). They are aligned with the curves obtained by the full MCF method up to the values of $i_{b}$ that are higher than the retrapping forces of the noiseless scenario [marked with the arrows at the bottom of Fig. 4(a)]. The mathematical explanation of this agreement is that for $i_{b}$ between the phase jump regime and the switching regime, the first two eigenvalues with the largest real parts are still sufficiently separated from the next ones. In addition, the comparison in Fig. 4(a) also shows that the multiple phase jumps smoothly change into the running phase as we enter the switching regime. Nevertheless, a different approach is needed to investigate the dynamics in this regime.
@@ III.2.2 Switching processes
The phase jumps within the double well play an important role even in the regime of larger bias forces, where the switching between running and locked solutions takes place. To show this and to evaluate the escape and retrapping rates between locked and running metastable states, we again introduce a simplified model. We divide the full probability distribution function into three well-separated regions and calculate the occupation probabilities for each of them. The sharp borders between the regions are defined by separatrices of the noiseless ($\Theta=0$) dissipative steady-state solution. In plain words, we determine the momentary regime of the particle by identifying the deterministic steady state in which the particle would end with its current position and velocity for the noiseless case.
@@
where $\Gamma_{eA}$ ($\Gamma_{eB}$) is the escape rate from potential well $A$ ($B$) to the metastable running state; $\Gamma_{rA}$ and $\Gamma_{rB}$ are the retrapping rates from metastable running state to the potential well $A$ or $B$, respectively, and $\Gamma_{AB}$ ($\Gamma_{BA}$) is the total rate of the phase jumps (of any length) from potential well $A$ to $B$ ($B$ to $A$) as it is illustrated in Fig. 4(b).
@@
Figure 6: Separatrices (black lines) shown for different dampings and bias forces. Panels can be used to illustrate complicated retrapping dynamics of the noiseless case. Any particle within the white region will keep running forever, particles within blue and red region will get trapped following paths bounded by the regions of the same color as their initial state (they can not cross separatrix). For example, a motionless particle placed initially close to the top of the higher potential maxima, but just right to the running-locked separatrix as marked by the green bullets ($v=0,$ $\varphi=\varphi^{b}+\delta$), will end in a minima marked by the green cross.
@@
In panel (a), where $\alpha=0.2$ and $\Theta=0.1$, the phase jump from minimum $A$ to minimum $B$ has a rate $\Gamma_{AB}$ (solid blue line) that is of the same order as the escape rate $\Gamma_{eA}$ (solid red line) and much higher than the escape rate $\Gamma_{eB}$ (dashed red line) at small bias. The rate of the opposite phase jump from $B$ to $A$ (blue dashed line) is negligible. In addition, the retrapping rate $\Gamma_{rB}$ (dashed black line) is almost two orders of magnitude larger than retrapping rate $\Gamma_{rA}$ (solid black line) in the entire plotted range. This means that the mean lifetime of a particle trapped in the locked state $A$ is negligible compared to the mean lifetime of the two other states. Consequently, in the steady state, the system exhibits bistable behavior reminiscent of a system governed by a simple single-harmonic potential. This is also evident in Fig. 7(b), where the stationary occupation probabilities are plotted. For low bias forces, the particle is predominantly trapped in the state $B$, while a sharp transition to the running state is observed near the Fano factor maximum (Fig. 4, blue curve). Throughout the range depicted in the figure, the locked state $A$ exhibits minimal influence on the system dynamics.
@@
Figure 7: (a),(c) Dependencies of the escape, retrapping, and phase jump rates [illustrated in Fig. 4(b)] on the bias force within the switching regime calculated using the simplified model (35). (b),(d) Related steady-state occupation probabilities. Two qualitatively different cases are presented: panels (a) and (b) show $\alpha=0.2$ where in the steady state the system behaves effectively as a single well, and panels (c) and (d) show $\alpha=0.02$ where the double-well character is evident.
@@
The situation is qualitatively different for $\alpha=0.02$. In this case, $\Gamma_{AB}$ and $\Gamma_{BA}$ are comparable and considerably smaller than those of the escape and retrapping processes within the range where switching occurs. However, this does not mean that phase jumps between the two locked states are irrelevant. As we discuss later in detail, it only means that their influence is apparent only on much longer timescales. At shorter ones, the phase dynamics is governed by the escape and retrapping processes. Due to the small $\alpha$, the rates of escape and retrapping processes are comparable ($\Gamma_{rA}\sim\Gamma_{rB}$ and $\Gamma_{eA}\sim\Gamma_{eB}$). Together, this leads to steady-state occupation probabilities where both locked states are relevant, as illustrated in Fig. 7(d). Interestingly, even with the low value of $\alpha=0.02$, there remains a significant difference between the occupation probabilities of states $A$ and $B$ in the steady state.
@@
Figure 8: Example of the time dependence of the occupation probabilities of the three metastable states for $i_{b}=0.2$ with initial condition $P_{R}=1$, $P_{A}=0$ and $P_{B}=0$. This models a sudden quench of very large bias force to bias force just above the retracing force of the noiseless case. Solid lines are results of the simplified model (35), dashed lines show full Langevin simulations for two different initial conditions.

### v2

@@ III.1 Strong damping
We start our analysis with the strong damping case. The main reason is that in this regime the dynamics is much simpler than for the intermediate and weak damping. Nevertheless, it is still far from trivial.
III.1.1 Fano factor and positional jumps
@@
Figure 4: (a) Examples of typical low-temperature Fano factor $F=S/(2\pi\langle v\rangle)$ for strong damping plotted as a function of bias force for different parameters $\alpha$. The inset illustrates how the $0.5$ Fano factor plateau transits into Fano plateau with value 1 as $\alpha$ increase. (b) A comparison of the $\alpha$-dependence of the single-positional-jump rates evaluated numerically (solid lines) with the analytical Kramers formula result for the overdamped case [83, 72] (dashed lines). The green bullets represent rates obtained via the simplified model where only $2\pi n$ PJs were taken into account.
@@
In Fig. 4(a) we show the Fano factor in the strongly damped regime represented by $\gamma=5$ for three values of $\alpha$ at temperature $\Theta=0.08$. The Fano factor exhibits a characteristic divergence close to the equilibrium ($i_{b}=0$) due to the finite thermal noise. For $\alpha=0$, it follows formula $F=\frac{1}{2}\coth(\pi i_{b}/2\Theta)$ [72, Eq. (28)] for small $i_{b}$ and $\alpha=0$ with a plateau at the Poissonian value of $F=0.5$ (note that we still normalize $F$ to $2\pi$). This is in agreement with the dominant contribution to the noise from the single PJ [illustrated in Fig. 1(c)] over a distance $d=\pi$ between the equivalent neighboring minima. The situation for the finite $\alpha$ is more complicated. The range in which single PJs are the dominant source of the overall noise is not marked by a clear plateau. Rather for $0<\alpha\ll 1$ [$\alpha=0.1$ in Fig. 4(a)] we observe a slight slope in the Fano factor. This reflects the fact that for finite $\alpha$ the prevailing velocity noise contribution consists of a nontrivial combination of the single positional jumps forward over the distance of $\varphi^{B}_{m}-\varphi^{A}_{m}=d$ and backward by $\varphi^{A}_{m+1}-\varphi^{B}_{m}=2\pi-d$ [see Fig. 1(c)] where $d$ depends on both $\alpha$ and $i_{b}$. As already stated, because $d\neq\pi$ we refer to these events as the fractional positional jumps.
@@
With increasing $\alpha$ the Fano factor increases smoothly in this region from $0.5$ signaling only PJ over the distance $\pi$ to $1$ suggesting a single PJs over $2\pi$ as shown in the inset of Fig. 4(a). Interestingly, for $\Theta=0.08$ the Fano factor approaches 1 even for values of $i_{b}$ that are still well below $i_{cA}$, therefore, in a regime where both minima still exist. For example, the critical current $i_{cA}$ is approximately $0.396$ at $\alpha=1$, yet its Fano factor follows the analytical result $F=\coth(\pi i_{b}/\Theta)$, valid for systems with only the first harmonic, even below this value. Nevertheless, this can be understood as a consequence of the large $\gamma$ and can be explained using the FCS method together with a simplified model of the elementary PJ processes.
@@
Because only jumps over a single (uneven) maxima are realized in the locked state of the strong-damping case the overall dynamics of this regime can be described using just four rates: $\Gamma^{A}_{1}$ and $\Gamma^{B}_{1}$ for the forward single jumps and $\Gamma^{A}_{-1}$ and $\Gamma^{B}_{-1}$ for the backward single jumps. Moreover, the backward rates can be neglected if the bias force is strong enough. The typical dependencies of $\Gamma^{A}_{1}$ and $\Gamma^{B}_{1}$ on $\alpha$ in this regime represented by the bias force $i_{b}=0.3$ and the damping $\gamma=5$ are plotted in Fig. 4(b) (solid lines). Both $\Gamma^{A}_{1}$ and $\Gamma^{B}_{1}$ closely follow the Kramers formula for escape across the adjacent barrier for overdamped case $\Gamma^{X}_{1}=\frac{1}{2\pi}\sqrt{|U^{\prime\prime}(\varphi^{x})|U^{\prime\prime}(\varphi^{X})}e^{-\Delta U_{X}/\Theta}$ [83, 72] (where $\Delta U_{X}=U(\varphi^{x})-U(\varphi^{X})$, and $x=a,b$; $X=A,B$) plotted with dashed lines of the respective colors. Note that because of the increasing difference $\Delta U_{B}-\Delta U_{A}$ the ratio
@@
increases exponentially with $\alpha$. Consequently, for high enough $\alpha$ and low enough $\Theta$ the average waiting time for the escape from minimum $A$ $\tau_{A\rightarrow B}=1/\Gamma^{A}_{1}$ is negligible compared to the waiting time for the escape from minimum $B$ ($\tau_{B\rightarrow A}=1/\Gamma^{B}_{1}$). Therefore, in the long term, every single positional jump $B_{m-1}\rightarrow A_{m}$ is immediately followed by a single positional jump $A_{m}\rightarrow B_{m}$. The combination of these two fractional PJs is effectively a complete $2\pi$ jump. This is shown in Fig. 4(b) by the green bullets that were obtained via a simplified model where only the $2\pi n$ jumps were considered [35]
@@
Their match with the $\Gamma^{B}_{1}$ rate for high enough $\alpha$ is consistent with the Fano-factor value of $1$ in the inset of Fig. 4(a). This has interesting physical consequences. If the temperature is low enough, then, because $\Theta$ is in the denominator of the exponent of Eq. (26), any strongly damped $\varphi$ junction or another equivalent system describable by a double harmonic potential with a finite bias $i_{b}$ will in the steady state resemble a simple single harmonic system. As such, it can be described by the analytical formulas derived for the single-harmonic potential.
@@ III.2 Intermediate and weak damping
The stochastic dynamics of the particle in the tilted double-well periodic potential in the regime of intermediate and weak damping is significantly richer than in the strong damping case. For example, for weak bias force, there are positional jumps over multiple maxima, and for stronger bias, a complicated switching between running and locked solutions is the prevailing source of the velocity noise. Even the retrapping of a particle from the running to locked regime has complex dynamics, as discussed below.
@@
We support this claim in Fig. 5(b) where an example of the stationary distribution function $W(v,\varphi)$ is plotted for the parameters (see figure description) close to the maximum of the peak. Here, the two distinct peaks at $v\approx 0$ centered around potential minima signal the two locked states. The continuous ridge that spreads above them is the running state. In this regime, the prevailing contribution to the overall velocity noise comes from the switching between these three well-separated metastable states.
When we lower $i_{b}$ to the regime in between the thermal divergence at $i_{b}\to 0$ and the switching maximum, multiple fractional positional jumps become the prevailing source of the velocity noise. We now analyze the regimes of positional jumps and switching separately and then show how they transition smoothly into each other.
Figure 5: (a) Three examples of a typical low-temperature Fano factor $F=S/(2\pi\langle v\rangle)$ dependence on the bias force for weak damping. The lines were calculated by the MCF method, and the bullets show the ME results from Eq. (25). Vertical arrows mark the related noiseless retrapping forces. (b) An illustration of the stationary distribution function $W(v,\varphi)$ in the switching regime calculated for $\gamma=0.2$, $\alpha=0.02$, $i_{b}=0.38$, and $\Theta=0.08$. The arrows illustrate the processes and their rates as used in Eq. (38). Namely, the red arrows represent escape from the locked states to the running one, the black arrows represent retrapping of the particle, and the blue arrows are the positional jumps between the locked states $A$ and $B$. The gray dashed curves show the separatrices between the running solution and the locked solution in the well $A$ (solid line) and/or well $B$ (dashed line), respectively.
III.2.1 Positional jumps
Figure 6: The rates of multiple fractional positional jumps (of order $n$) for different values of the bias force $i_{b}$ and ratio $\alpha$. The inset in (a) is a verification of the detailed balance condition. The blue columns describe jumps starting at minimum $A$, red at minimum $B$ (a small horizontal shift is introduced for visibility), and the green columns in panel (c) show the rates obtained via the simplified model where only $2\pi n$ PJs were considered [analogously to Fig. 4(b)].
@@ III.2.1 Positional jumps
The profile of the rates for potentials with finite $\alpha$ is more complicated. The $n$ dependencies of the PJ rates differ qualitatively between jumps starting in different minima, as well as between odd and even $n$’s. We show this in Figs. 6(b)-(c). The rates $\Gamma^{A}_{n}$ and $\Gamma^{B}_{n}$ differ by several orders of magnitude for odd $n$’s [see $\alpha=0.2$ case in Fig. 6(c)], but are comparable for even $n$’s. This can be rationalized by analyzing the “trajectories” of the particular PJs following the illustrations in Fig. 1(b)(c).
The jumps $A_{m}\rightarrow A_{m+n/2}$ and $B_{m}\rightarrow B_{m+n/2}$, where $n$ is an even integer, are indeed comparable. Here, the particle had to overcome the same number of maxima $a$ as well as $b$, namely $n/2$. In the idealized case, they traveled the same distance $\pi n$. Therefore, also the related rates are equivalent. However, this is not true for odd jumps representing fractional PJs between minima of different kinds. In the case of an odd $n$ the particle overcomes $(n+1)/2$ of the lower maxima but only $(n-1)/2$ of the higher ones during a $\Gamma^{A}_{n}$ process. The opposite is true for the $\Gamma^{B}_{n}$ process. Furthermore, travel distances differ by $2\pi-2d$. Consequently, the “odd” rates differ significantly ($\Gamma^{A}_{n}\gg\Gamma^{B}_{n}$ for positive $n$).
@@
Before moving to the switching regime, it is worth stressing that the FCS method works well up to surprisingly high Fano factors ($F\sim 10^{2}$). This is shown in Fig. 5(a) where the bullets represent the Fano factor obtained directly from the rates by Eq. (25). They are aligned with the curves obtained by the full MCF method up to the values of $i_{b}$ that are higher than the retrapping forces of the noiseless scenario [marked with the arrows at the bottom of Fig. 5(a)]. The mathematical explanation of this agreement is that for $i_{b}$ between the positional-jump regime and the switching regime, the first two eigenvalues with the largest real parts are still sufficiently separated from the next ones. In addition, the comparison in Fig. 5(a) also shows that the multiple positional jumps smoothly change into the running state as we enter the switching regime. Nevertheless, a different approach is needed to investigate the dynamics in this regime.
@@ III.2.2 Switching processes
The positional jumps within the double well play an important role even in the regime of larger bias forces, where the switching between running and locked solutions takes place. To show this and to evaluate the escape and retrapping rates between locked and running metastable states, we again introduce a simplified model. We divide the full probability distribution function into three well-separated regions and calculate the occupation probabilities for each of them. The sharp borders between the regions are defined by separatrices of the noiseless ($\Theta=0$) dissipative steady-state solution. In plain words, we determine the momentary regime of the particle by identifying the deterministic steady state in which the particle would end with its current position and velocity for the noiseless case.
@@
where $\Gamma_{eA}$ ($\Gamma_{eB}$) is the escape rate from potential well $A$ ($B$) to the metastable running state; $\Gamma_{rA}$ and $\Gamma_{rB}$ are the retrapping rates from metastable running state to the potential well $A$ or $B$, respectively, and $\Gamma_{AB}$ ($\Gamma_{BA}$) is the total rate of the positional jumps (of any length) from potential well $A$ to $B$ ($B$ to $A$) as illustrated in Fig. 5(b).
@@
In panel (a), where $\alpha=0.2$ and $\Theta=0.1$, the positional jump from minimum $A$ to minimum $B$ has a rate $\Gamma_{AB}$ (solid blue line) that is of the same order as the escape rate $\Gamma_{eA}$ (solid red line) and much higher than the escape rate $\Gamma_{eB}$ (dashed red line) at small bias. The rate of the opposite positional jump from $B$ to $A$ (blue dashed line) is negligible. In addition, the retrapping rate $\Gamma_{rB}$ (dashed black line) is almost two orders of magnitude larger than retrapping rate $\Gamma_{rA}$ (solid black line) in the entire plotted range. This means that the mean lifetime of a particle trapped in the locked state $A$ is negligible compared to the mean lifetime of the two other states. Consequently, in the steady state, the system exhibits bistable behavior reminiscent of a system governed by a simple single-harmonic potential. This is also evident in Fig. 7(b), where the stationary occupation probabilities are plotted. For low bias forces, the particle is predominantly trapped in the state $B$, while a sharp transition to the running state is observed near the Fano factor maximum (Fig. 5, blue curve). Throughout the range depicted in the figure, the locked state $A$ exhibits minimal influence on the system dynamics.
@@
Figure 7: (a),(c) Dependencies of the escape, retrapping, and positional jump rates [illustrated in Fig. 5(b)] on the bias force within the switching regime calculated using the simplified model (34). (b),(d) Related steady-state occupation probabilities. Two qualitatively different cases are presented: panels (a) and (b) show $\alpha=0.2$ where in the steady state the system behaves effectively as a single well, and panels (c) and (d) show $\alpha=0.02$ where the double-well character is evident.
Figure 8: Comparison of the factor factor $F=S/(2\pi\langle v\rangle)$ dependencies on the bias force where the black lines were calculated by the MCF method, and the red circles show the ME results from Eq. (40).
@@
The situation is qualitatively different for $\alpha=0.02$. In this case, $\Gamma_{AB}$ and $\Gamma_{BA}$ are comparable and considerably smaller than those of the escape and retrapping processes within the range where switching occurs. However, this does not mean that positional jumps between the two locked states are irrelevant. As we discuss later in detail, it only means that their influence is apparent only on much longer timescales. At shorter ones, the dynamics is governed by the escape and retrapping processes. Due to the small $\alpha$, the rates of escape and retrapping processes are comparable ($\Gamma_{rA}\sim\Gamma_{rB}$ and $\Gamma_{eA}\sim\Gamma_{eB}$). Together, this leads to steady-state occupation probabilities where both locked states are relevant, as illustrated in Fig. 7(d). Interestingly, even with the low value of $\alpha=0.02$, there remains a significant difference between the occupation probabilities of states $A$ and $B$ in the steady state.
The results of the simplified model (34) can also be used to test whether the peak in the Fano factor actually arises as a result of switching between the running and locked states. In the vicinity of the peak where $F\gg 1$ we can neglect the noise contributions inherent to particular metastable states [84, 35] and calculate $F$ from the average velocity (obtained here from MCF) and rates and probabilities from the ME (34) as [35]
$F_{\mathrm{ME}}=\frac{\langle v\rangle\left(v_{r}-\langle v\rangle\right)}{\pi v_{r}\Gamma_{e}},$ (40)
where we approximate the velocity in the running state by $v_{r}=i_{b}/\gamma$. The total escape rate $\Gamma_{e}$ is calculated as $\Gamma_{e}=\bar{P}_{A}\Gamma_{eA}+\bar{P}_{B}\Gamma_{eB}$ where $\bar{P}_{A,B}=P_{A,B}/(P_{A}+P_{B})$ reflect the probabilities with which the already locked particle is placed in minima $A$ or $B$. We show in Fig. 8 a comparison of the Fano factors obtained from Eq. (40) (red circles) with the full MCF solution (black lines) for the same parameters as used in Fig. 7. The Fano factors are in a very good agreement confirming both the origin of the peak from the switching processes and the validity of the simplified model (34).
@@
Figure 9: Example of the time dependence of the occupation probabilities of the three metastable states for $i_{b}=0.2$ with initial condition $P_{R}=1$, $P_{A}=0$ and $P_{B}=0$. This models a sudden quench of very large bias force to bias force just above the retrapping force of the noiseless case. Solid lines are results of the simplified model (34), dashed lines show full Langevin simulations for two different initial conditions.